# Redistricting Reforms Reduce Gerrymandering by Constraining Partisan Actors

### v1

@@ 7 Conclusion
In this paper, we propose a methodology for estimating causal effects of complex institutional reforms and apply it to ascertain the impacts of redistricting reforms in the United States. Although redistricting reforms differ across states in important procedural details, we show how to obtain a theoretically informed univariate summary measure of such reforms through a game theoretic model. We then combine this approach with a standard causal inference research design strategy to draw a credible causal inference even with a limited sample size. We find that redistricting reforms reduce the partisan bias of enacted plans by constraining the leeway of partisan actors.
Our methodological approach enables us to go beyond the estimation of causal effects. Specifically, we conduct a counterfactual policy analysis to infer the consequences of different nationwide redistricting reforms. We find that instituting redistricting commissions generally reduces the current Republican advantage, but the details of commission structure matter. For example, Michigan-style reform would yield a much greater pro-Democratic effect than types of redistricting commissions adopted in Ohio and New York. While we apply this approach to three commission reforms, this model can be applied to other reforms of interest.
The key idea behind our methodology is also applicable to studies of other complex institutions. For example, most scholars analyze democratic institutions by using the Polity score, which is an 21-point scale index based on six “component variables” concerning executive recruitment, executive constraints, and political participation. Instead of simply summing these scores to obtain the overall Polity score, our proposed methodology suggests that researchers consider developing a more theoretically informed measure of democracy, based directly on their substantive questions of interest. We believe that such an approach preserves critical aspects of institutional complexity while improving the credibility of causal analysis.
@@
In Ohio 2020, if the legislature fails to redistrict by bipartisan majorities, the work falls to a commission that must approve a map by a bipartisan vote. The commission is composed of appointees by state executive officers and legislative parties. If the commission cannot do so, it can pass a map by a majority vote without bipartisan support but the map only lasts two election cycles instead. This is encoded as a draw by a split state legislature with the first stalemate broken by a nonpartisan commission and the second stalemate broken by Republicans, since Republicans control all state executive offices and have a functional majority on the commission.
@@
In Virginia in 2010, the state legislature initially deadlocked, but drew a map in time after state legislative elections in 2011. Thus the stalemate procedure for 2010 was coded as unclear.
@@
The game tree is specified in Figure 2 and is described in Section 3.2. Here we detail the parametric specification of the moves by nature, starting with the post-enactment court review process.
@@
The prior on chal_poss_conf is and the prior on chal_poss_maybe is . The effect of this setup is to follow the data closely but allow for some uncertainty in the possibility of court review, rather than, e.g., precluding any possibility of court challenge in a state because it was coded as “No.” State courts may change their interpretations of state laws and constitutions in any given case.
@@
When veto control is coded as as NA (other than in the case of a governor with overridable veto) or a court has a veto, the veto is exercised with probability
@@
Split-control veto players are assumed to never exercise their veto, as are governors that can be overridden.
@@
Map drawers with split partisan control are assumed to stalemate with probability controlled by parameter stale_split_prob, which has a prior. When they don’t stalemate, split-control map drawers are assumed to propose a map with the same partisan bias as one drawn by a nonpartisan map drawer.
@@
The veto_nonp_shift parameter controls the amount of the shift and receives a prior.

### v2

@@ 7 Conclusion
In this paper, we propose a methodology for estimating causal effects of complex institutional reforms and apply it to study the impacts of redistricting reforms in the United States. Although redistricting reforms differ across states in important procedural details, we show how to obtain a theoretically informed univariate summary measure of such reforms through a game-theoretic model. We then combine the results of this formal model with a standard causal inference research design strategy to obtain credible causal effect estimates even with a limited sample size. We find that redistricting reforms reduce the partisan bias of enacted plans by constraining the leeway of partisan actors.
Our methodological approach enables us to go beyond the estimation of causal effects. Specifically, we conduct a counterfactual policy analysis to infer the consequences of adopting a series of different redistricting reforms nationwide. We find that adopting redistricting commissions generally reduces the current Republican advantage, but commissions are more effective at reducing bias when they place greater restrictions on partisan actors. For example, a Michigan-style reform (which combines a nonpartisan commission with no partisan veto points) would yield a much greater pro-Democratic effect than commissions adopted in Ohio and New York (which maintain the potential for some partisan control). While we apply this approach to three commission reforms, the same model can be applied to other reforms of interest.
The key idea behind our methodology is also applicable to studies of other complex institutions. For example, most scholars analyze democratic institutions by using the Polity score, which is a 21-point scale index based on six “component variables” concerning executive recruitment, executive constraints, and political participation. Instead of summing these scores to obtain the overall Polity score, our proposed methodology suggests that researchers consider developing a more theoretically informed measure of democracy, based directly on their substantive questions of interest. We believe that such an approach preserves critical aspects of institutional complexity while improving the credibility of causal analysis.
@@
In Ohio 2020, if the legislature fails to redistrict by bipartisan majorities, the work falls to a commission that must approve a map by a bipartisan vote. The commission is composed of appointees by state executive officers and legislative parties. If the commission cannot do so, it can pass a map by a majority vote without bipartisan support, but the map only lasts two election cycles instead. This is encoded as a draw by a split state legislature with the first stalemate broken by a nonpartisan commission and the second stalemate broken by Republicans, since Republicans control all state executive offices and have a functional majority on the commission.
@@
In Virginia in 2010, the state legislature initially deadlocked, but drew a map in time after state legislative elections in 2011. Thus, the stalemate procedure for 2010 was coded as unclear.
![[Uncaptioned image]](x8.png)
@@
The game tree is specified in Figure 2 and is described in Section 3.2. Here we detail the parametric specification of the moves by nature, starting with the post-enactment court review process. The parameters in the model all receive probabilistic priors, rather than being fixed to specific values. These priors place most of their mass across a broad range of plausible values, based on our experience and observation of past redistricting processes.
@@
The prior on chal_poss_conf is and the prior on chal_poss_maybe is . The effect of this setup is to follow the treatment coding closely but allow for some uncertainty in the possibility of court review, rather than, e.g., precluding any possibility of court challenge in a state because it was coded as “No.” State courts may change their interpretations of state laws and constitutions in any given case.
@@
When veto control is coded as NA (other than in the case of a governor with overridable veto) or a court has a veto, the veto is exercised with probability
@@
[…]
Map drawers with split partisan control are assumed to stalemate with probability controlled by parameter stale_split_prob, which has a prior. When they don’t stalemate, split-control map drawers are assumed to propose a map with the same partisan bias as one drawn by a nonpartisan map drawer. While we suspect this probability could be set to 1 without changing the results, a probability of compromise reflects a history of compromise plans in state legislatures.
@@
The veto_nonp_shift parameter controls the amount of the shift and receives a prior. This mechanism reflects real-world patterns of repeated map-drawing, such as when the (Republican-controlled) Iowa legislature rejected the proposed plan from the nonpartisan advisory commission, then adopted a second proposal from the same commission that favored Republicans. more.
[…]
In Alabama, the legislature draws congressional districts, subject to a governor’s veto, and there is no mechanism for state court review. Thus, Alabama’s process would be described by the “Drawer” and “Veto 1” steps only in Round 1 of the Figure 2 process.
The first move belongs to the Republican party, which controls the state legislature. It has to pick the amount of partisan bias in the plan that it adopts. If this plan is ultimately the final enacted plan, the Republicans receive utility and the Democrats receive utility . The second move also belongs to the Republican party, which controls the governorship. The party must decide whether or not to veto the adopted plan; if it does, there is a second round of drawing and vetoing by the legislature and governor (controlled by Republicans).
[…]